OpenStudy (anonymous):

Solve the word problem. Jackie mowed a lawn today to earn some extra money. She spent $1.90 on fuel and $4.65 on lunch. She charged $12.50 to mow the lawn. What was her profit for mowing the lawn? A. $5.95 B. $9.75 C. $15.25 D. $19.05

OpenStudy (anonymous):

@mossyfish

OpenStudy (johnweldon1993):

Well she MADE 12.50 she SPENT 1.90 + 4.65 = 6.55 So she PROFITS 12.50 - 6.55 = 5.95

OpenStudy (anonymous):

a 5.95

OpenStudy (anonymous):

Let's add her expenses. \[$1.90+$4.65=e\] What does e equal? Once you figure that out, simply plug in your result into this equation: \[12.50-e=profit\]